OpenStudy (anonymous):

What is the solution of the system? Use substitution. 3x + 2y = 7 y = –3x + 11 A. (6, -3) B. (6, -7) C. (-4,19/2) D. (5,-4)

OpenStudy (anonymous):

@Qwertty123 @alexandrea.lane

jagr2713 (jagr2713):

Hey, what do you think we do first?

OpenStudy (anonymous):

factor?

jagr2713 (jagr2713):

No, when substituting we are trying to get x/y alone and plug it back in to get x/y alone. So basically we are trying to get 2 points to graph.

OpenStudy (anonymous):

ok. how do i do that

jagr2713 (jagr2713):

Well, they already set it equals to Y for us so all we do is plug Y in into the first equation. What would we do?

OpenStudy (anonymous):

subtract 11 and 3x?

OpenStudy (anonymous):

i have no idea

OpenStudy (qwertty123):

3x + 2y =7 Y = -3x + 11 adding them:

OpenStudy (qwertty123):

2y+11=y+7

jagr2713 (jagr2713):

Here is a better explanation "The method of solving "by substitution" works by solving one of the equations (you choose which one) for one of the variables (you choose which one), and then plugging this back into the other equation, "substituting" for the chosen variable and solving for the other. Then you back-solve for the first variable."

jagr2713 (jagr2713):

3x + 2y = 7 y = –3x + 11 Plug it in for Y and get 3x+2(-3x+11)=7

jagr2713 (jagr2713):

So, what do we do first in the equation?

OpenStudy (anonymous):

so distribute?

jagr2713 (jagr2713):

Yes, what do we get?

OpenStudy (anonymous):

-6x+22???

jagr2713 (jagr2713):

Correct so we have 3x-6x+22=7 What do we do next?

OpenStudy (anonymous):

-3x+22=7?

OpenStudy (anonymous):

-22?

jagr2713 (jagr2713):

-22? what do you mean?

OpenStudy (anonymous):

-3x=-15 divide by -3= 5?

jagr2713 (jagr2713):

What do we get?

OpenStudy (anonymous):

5

jagr2713 (jagr2713):

COrrect :D what do we do next?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

5,-4???

jagr2713 (jagr2713):

Now, we plug 5 into the equation to get Y y = –3(5) + 11

OpenStudy (anonymous):

ok

jagr2713 (jagr2713):

What do we get?

OpenStudy (anonymous):

-4

jagr2713 (jagr2713):

Correct :D Great job so the answer is ?

OpenStudy (anonymous):

5,-4 is answer

OpenStudy (anonymous):

thanks

jagr2713 (jagr2713):

Correct